# 3cm3cm3cm3cm

### v1

@@ 1. Introduction
Notations. Throughout the paper, the symbol ${\mathbb{Z}}$ (resp., ${\mathbb{N}}$ and ${\mathbb{Z}}_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol ${\mathbb{C}}$ for the field of complex numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. All vector spaces, algebras, tensor products, etc., are over ${\mathbb{C}}$. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
@@ 2. Preliminaries
In this section, we set up our notations for the general linear Lie (super)algebras and their related concepts. We review several notions needed in this paper, including current algebras, the Feigin–Frenkel centers and the Bethe algebras.
2.1. The general linear Lie (super)algebras
@@
[…]
@@
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of ${\mathbb{C}}$-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called a general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$. The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
$[E^{i}_{j},E^{k}_{\ell}]=\delta_{j,k}E^{i}_{\ell}-(-1)^{(|i|+|j|)(|k|+|\ell|)}\delta_{i,\ell}E^{k}_{j}\qquad\mbox{for $i,j,k,\ell\in\mathbb{I}$.}$
@@
Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$. The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$, and the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ is denoted by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$, where the parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$. Note that the Borel subalgebra $\mathfrak{b}_{{p+m|q+n}}$ is not the standard one unless some of $p,q,m,n$ are set to 0.
For $m=d$ and $p=q=n=0$, we write
${\mathfrak{gl}}_{d}={\mathfrak{gl}}_{d|0},\quad{\mathfrak{b}}_{d}={\mathfrak{b}}_{d|0},\quad{\mathfrak{h}}_{d}={\mathfrak{h}}_{d|0},$
and
$\hskip 28.45274pte_{ij}=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$
2.2. Current algebras and evaluation modules
For any Lie (super)algebra ${\mathfrak{g}}$ and an even variable $t$, let ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ denote the loop algebra of ${\mathfrak{g}}$, which is a Lie superalgebra with commutation relations
@@ 2.2. Current algebras and evaluation modules
The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ is a subalgebra of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way. We identify the Lie (super)algebra ${\mathfrak{g}}$ with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$. Hereafter $U(\mathcal{G})$ denotes the universal enveloping algebra of $\mathcal{G}$ for any Lie (super)algebra $\mathcal{G}$.
@@
Any ${\mathfrak{g}}$-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, which is called the evaluation module and denoted by $M(\alpha)$. For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
@@
Here every rational function in $z$ represents its power series expression at $\infty$.
@@
$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$
@@
2.3. The Feigin–Frenkel center
The Lie algebra ${\mathfrak{gl}}_{d}$ admits a nondegenerate invariant symmetric bilinear form
$(A,B)={\rm tr}(AB)-\frac{1}{d}{\rm tr}(A){\rm tr}(B),\quad\mbox{for $A,B\in{\mathfrak{gl}}_{d}$},$
where ${\rm tr}$ denotes the trace. The affine Lie superalgebra
$\widehat{{\mathfrak{gl}}_{d}}:={\mathfrak{gl}}_{d}[t,t^{-1}]\oplus{\mathbb{C}}K,$
where $K$ is even and central, is a Lie superalgebra with commutation relations
$\big{[}A[r],B[s]\big{]}=[A,B][r+s]+r\delta_{r+s,0}(A,B)K$
for $A,B\in{\mathfrak{gl}}_{d}$ and $r,s\in{\mathbb{Z}}$.
The vacuum module $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ at the critical level is
$V_{{\rm crit}}({\mathfrak{gl}}_{d}):=U(\widehat{{\mathfrak{gl}}_{d}})/I,$
where $I$ is the left ideal of $U(\widehat{{\mathfrak{gl}}_{d}})$ generated by ${\mathfrak{gl}}_{d}[t]$ and $K+d$. It possesses a unique vertex algebra structure, called the universal affine vertex algebra associated to ${\mathfrak{gl}}_{d}$ at the critical level; see [FBZ, Theorem 4.4.1]. The center of the vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{d})$ is given by
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d}):=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{gl}}_{d})\;\middle|\;{\mathfrak{gl}}_{d}[t]v=0\!\,\right\},$
called the Feigin–Frenkel center. Any element of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is called a Segal–Sugawara vector.
We also have the universal affine vertex algebra $V_{{\rm crit}}({\mathfrak{gl}}_{p+m|q+n})$ at the critical level and the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for ${\mathfrak{gl}}_{p+m|q+n}$. We skip details here; see [MR, Section 1.3] (and also [ChL2, Section 3.1]).
Let ${\mathfrak{g}}$ be a general linear Lie (super)algebra. The algebra $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is equipped with the (even) derivation ${\rm T}:=-d/dt$ defined by
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$
for $A\in{\mathfrak{g}}$ and $r\in{\mathbb{N}}$. Note that $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ is isomorphic to $V_{{\rm crit}}({\mathfrak{g}})$ as superspaces by the Poincaré–Birkhoff–Witt theorem, and that the derivation ${\rm T}$ corresponds to the translator operator on $V_{{\rm crit}}({\mathfrak{g}})$. It is known that ${\mathfrak{z}}(\widehat{{\mathfrak{g}}})$ can be viewed as a commutative subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ and is ${\rm T}$-invariant. We refer the reader to [Fr, FBZ, Mo, MR] for further details.
@@ 2.4. Bethe algebras
For $\mu\in{\mathfrak{gl}}_{d}^{*}$, there is an algebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])\longrightarrow U({\mathfrak{gl}}_{d}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
@@
$\hskip 28.45274pt\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\quad\mbox{for $A\in{\mathfrak{gl}}_{d}\,$ and $\,r\in{\mathbb{N}}$.}$ (2.3)
The Bethe algebra $\mathcal{B}_{d}^{\mu}$ of ${\mathfrak{gl}}_{d}$ with respect to $\mu$ is defined to be the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $\Psi^{\mu}(S)$ for $S\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$ is commutative, so is $\mathcal{B}_{d}^{\mu}$ ([FFTL, Ry]).
For any $\mathcal{B}_{d}^{\mu}$-module $V$, let $(\mathcal{B}_{d}^{\mu})_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{d}^{\mu}$ in the endomorphism algebra ${\rm End}(V)$ of $V$. We call $(\mathcal{B}_{d}^{\mu})_{V}$ the Bethe algebra of $V$ with respect to $\mu$. For $\alpha\in{\mathbb{C}}$, any ${\mathfrak{gl}}_{d}$-module $M$ is the evaluation module $M(\alpha)$ over ${\mathfrak{gl}}_{d}[t]$, and so $\mathcal{B}_{d}^{\mu}$ acts on $M(\alpha)$. If the centralizer ${\mathfrak{gl}}_{d}^{\mu}$ of $\mu$ in ${\mathfrak{gl}}_{d}$ contains the Cartan subalgebra ${\mathfrak{h}}_{d}$, then $\mathcal{B}_{d}^{\mu}$ commutes with ${\mathfrak{h}}_{d}$ and hence acts on any $\gamma$-weight space $M(\alpha)_{\gamma}$ of $M(\alpha)$, where $\gamma\in{\mathfrak{h}}_{d}^{*}$ ([Ry, Proposition 4]).
Let
${\boldsymbol{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\}$
be the configuration space of $\ell$ distinct points on ${\mathbb{C}}^{\ell}$, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules. The Bethe algebra $\mathcal{B}_{d}^{\mu}\subseteq U({\mathfrak{gl}}_{d}[t])$ acts on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ (for the notation, see (2.2)) via the evaluation map $\underline{{\rm ev}}_{\bf z}$.
An element $\mu\in{\mathfrak{gl}}_{d}^{*}$ is called regular if the dimension of ${\mathfrak{gl}}_{d}^{\mu}$ is $d$. The following remarkable results for $\mathcal{B}_{d}^{\mu}$, due to Feigin, Frenkel, and Rybnikov, will be used in this paper.
Theorem 2.1 (cf. [FFRy, Corollary 5]).
@@
For any regular $\mu\in{\mathfrak{gl}}_{d}^{*}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{d}^{\mu}$-module.
Theorem 2.2 (cf. [FFRy, Corollary 6]).

### v2

@@ 1.3. Outline
Notations. Throughout the paper, the symbol $\mathbb{Z}$} (resp., $\mathbb{N}$} and $\mathbb{Z}$$_{+}$) stands for the set of all (resp., positive and non-negative) integers, the symbol $\mathbb{C}$} (resp., $\mathbb{R}$}) for the field of complex (resp., real) numbers, and the symbol ${\mathbb{Z}}_{2}:=\{\bar{0},\bar{1}\}$ for the field of integers modulo 2. Unless otherwise stated, all vector spaces, algebras, tensor products, etc., are over $\mathbb{C}$}. We fix $d\in{\mathbb{N}}$ and $p,q,m,n\in{\mathbb{Z}}_{+}$.
@@ 2. Preliminaries
In this section, we review the classical Lie (super)algebras of types $\mathfrak{a,c,d}$, evaluation modules, and pseudo-differential operators.
2.1. Classical Lie (super)algebras
2.1.1. The general linear Lie (super)algebra
@@
Let $\{e_{i}\,|\,i\in\mathbb{I}\}$ be a basis for the superspace ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ such that $\{e_{i}\,|\,1\leq i\leq p+q\}$ and $\{e_{p+q+i}\,|\,1\leq i\leq m+n\}$ are respectively the standard homogeneous bases for ${\mathbb{C}}^{p|q}$ and ${\mathbb{C}}^{m|n}$. In other words, the parity of $e_{i}$ is given by $|e_{i}|=|i|$ for $i\in\mathbb{I}$.
@@
where $\delta$ is the Kronecker delta. The parity of $E^{i}_{j}$ is given by $|E^{i}_{j}|=|i|+|j|$. The superspace of $\mathbb{C}$}-linear endomorphisms on ${\mathbb{C}}^{p|q}\oplus{\mathbb{C}}^{m|n}$ has a natural structure of a Lie superalgebra, called the general linear Lie (super)algebra and denoted by ${\mathfrak{gl}}_{p+m|q+n}$. The commutation relations of ${\mathfrak{gl}}_{p+m|q+n}$ are given by
$E^{i}_{j},E^{k}_{l}]=\delta_{j,k}E^{i}_{l}-(-1)^{(|i|+|j|)(|k|+|l|)}\delta_{i,l}E^{k}_{j}\qquad\mbox{for $i,j,k,l\in\mathbb{I}$.}$
@@
Let $\mathfrak{b}_{{p+m|q+n}}=\bigoplus_{\begin{subarray}{c}i,j\in\mathbb{I},i\leq j\end{subarray}}{\mathbb{C}}E^{i}_{j}$ be a Borel subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$. The corresponding Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ has a basis $\{E^{i}_{i}\,|\,i\in\mathbb{I}\}$. We denote the dual basis in ${\mathfrak{h}}_{{p+m|q+n}}^{*}$ by $\{\epsilon_{i}\,|\,i\in\mathbb{I}\}$. The parity of $\epsilon_{i}$ is given by $|\epsilon_{i}|=|i|$.
For $m=d$ and $p=q=n=0$, we define
${\mathfrak{gl}}_{d}:={\mathfrak{gl}}_{d|0}\qquad\text{and}\qquad e^{\mathfrak{a}}_{ij}:=e_{ij}:=E^{i}_{j}\qquad\mbox{for $i,j=1,\ldots,d$.}$ (2.1)
2.1.2. The symplectic Lie algebra
The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate skew-symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
$\left[\begin{matrix}0&J_{d}\\ -J_{d}&0\end{matrix}\right]$
is called the symplectic Lie algebra and is denoted by $\mathfrak{sp}_{2d}$. Hereafter, $J_{d}$ is the $d\times d$ reversion matrix defined by $(J_{d})_{ij}=1$ if $i+j=d+1$ and $(J_{d})_{ij}=0$ otherwise.
The Lie algebra $\mathfrak{sp}_{2d}$ is simple and is spanned by
$e^{\mathfrak{c}}_{ij}:=e_{ij}-\varepsilon_{i}\varepsilon_{j}e_{j^{\prime}i^{\prime}}$ (2.2)
for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and $\varepsilon_{i}:=1$ if $1\leq i\leq d$ and $\varepsilon_{i}:=-1$ if $d+1\leq i\leq 2d$.
2.1.3. The orthogonal Lie algebra
The subalgebra of ${\mathfrak{gl}}_{2d}$ that preserves the nondegenerate symmetric bilinear form on ${\mathbb{C}}^{2d}$ corresponding to the matrix
$\left[\begin{matrix}0&J_{d}\\ J_{d}&0\end{matrix}\right]$
is called the orthogonal Lie algebra and is denoted by $\mathfrak{so}_{2d}$. It is spanned by
$e^{\mathfrak{d}}_{ij}:=e_{ij}-e_{j^{\prime}i^{\prime}}$ (2.3)
for $i,j=1,\ldots,2d$, where $i^{\prime}:=2d-i+1$, and is a simple Lie algebra if $d\geq 3$.
2.1.4. Triangular decompositions
The general linear Lie (super)algebras ${\mathfrak{gl}}_{p+m|q+n}$ and ${\mathfrak{gl}}_{d}$ are called the classical Lie (super)algebras of types $\mathfrak{a}$ while the Lie algebras $\mathfrak{sp}_{2d}$ and $\mathfrak{so}_{2d}$ are called the classical Lie algebras of types $\mathfrak{c}$ and $\mathfrak{d}$, respectively. Let
${\mathfrak{g}}^{\mathfrak{a}}={\mathfrak{gl}}_{d},\qquad{\mathfrak{g}}^{\mathfrak{c}}=\mathfrak{sp}_{2d},\qquad{\mathfrak{g}}^{\mathfrak{d}}=\mathfrak{so}_{2d}.$
For ${\mathfrak{x}}=\mathfrak{a,c,d}$, the classical Lie algebra ${\mathfrak{g}}^{\mathfrak{x}}$ has the triangular decomposition
${\mathfrak{g}}^{\mathfrak{x}}=\mathfrak{n}_{-}^{\mathfrak{x}}\oplus{\mathfrak{h}}^{\mathfrak{x}}\oplus\mathfrak{n}_{+}^{\mathfrak{x}},$ (2.4)
where ${\mathfrak{h}}^{\mathfrak{x}}:=\bigoplus_{i=1}^{d}{\mathbb{C}}e^{\mathfrak{x}}_{ii}$ is the standard Cartan subalgebra of ${\mathfrak{g}}^{\mathfrak{x}}$, and $\mathfrak{n}_{+}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}e^{\mathfrak{x}}_{ij}$ and $\mathfrak{n}_{-}^{\mathfrak{x}}:=\sum_{\begin{subarray}{c}i<j\end{subarray}}{\mathbb{C}}e^{\mathfrak{x}}_{ji}$ are nilpotent subalgebras of ${\mathfrak{g}}^{\mathfrak{x}}$. We will suppress the superscript $\mathfrak{x}$} if there is no danger of confusion.
2.2. Evaluation modules
For any Lie superalgebra $\mathfrak{g}$}, we denote by $U({\mathfrak{g}})$ the universal enveloping algebra of $\mathfrak{g}$}. For any even variable $t$, the loop algebra ${\mathfrak{g}}[t,t^{-1}]:={\mathfrak{g}}\otimes{\mathbb{C}}[t,t^{-1}]$ is defined to be the Lie superalgebra with commutation relations
@@ 2.2. Evaluation modules
The current algebra ${\mathfrak{g}}[t]:={\mathfrak{g}}\otimes{\mathbb{C}}[t]$ and the superalgebra $t^{-1}{\mathfrak{g}}[t^{-1}]:={\mathfrak{g}}\otimes t^{-1}{\mathbb{C}}[t^{-1}]$ are subalgebras of ${\mathfrak{g}}[t,t^{-1}]$ defined in an obvious way. We identify $\mathfrak{g}$} with the subalgebra ${\mathfrak{g}}\otimes 1$ of constant polynomials in ${\mathfrak{g}}[t]$ and hence $U({\mathfrak{g}})\subseteq U({\mathfrak{g}}[t])$.
@@
Any $\mathfrak{g}$}-module $M$ is a ${\mathfrak{g}}[t]$-module via ${\rm ev}_{\alpha}$, called the evaluation module and denoted by $M(\alpha)$. For any $A\in{\mathfrak{g}}$ and any even variable $z$, we set
@@
Here each rational function in $z$ represents its power series expression at $\infty$.
@@
$\underline{{\rm ev}}_{\bf z}:=({\rm ev}_{z_{1}}\otimes\ldots\otimes{\rm ev}_{z_{\ell}})\circ\Delta^{(\ell-1)}:U({\mathfrak{g}}[t])\longrightarrow U({\mathfrak{g}})^{\otimes{\ell}},$ (2.5)
@@
2.3. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra and $z$ an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ (resp., $\mathcal{A}(\!(z)\!)$) the superalgebra of formal power series (resp., formal Laurent series) in $z$ with coefficients in $\mathcal{A}$. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
$\sum_{j=-\infty}^{s}\sum_{i=-\infty}^{r}a_{ij}z^{i}\partial_{z}^{j},$
where $r,s\in{\mathbb{Z}}$ and $a_{ij}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{i}z^{j}=\sum_{k=0}^{\infty}\binom{i}{k}\binom{j}{k}k!\,z^{j-k}\,\partial_{z}^{i-k},\quad\mbox{for $i,j\in{\mathbb{Z}}$.}$ (2.8)
Here, e.g., ${\binom{i}{k}:=\frac{i(i-1)\ldots(i-k+1)}{k!}}$. The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
$\omega(az^{i}\partial_{z}^{j})=\omega(a)z^{j}\partial_{z}^{i}$ (2.9)
for $a\in\mathcal{A}$ and $i,j\in{\mathbb{Z}}$. It is straightforward to prove the following.
Proposition 2.1.
3. Bethe algebras for classical Lie algebras
In this section, we examine the behavior of the Feigin–Frenkel center and the Bethe algebra for the classical Lie algebra $\mathfrak{g}$} of type $\mathfrak{a,c,d}$ under a *-structure. Our primary purpose is to study the diagonalization of the Bethe algebra for $\mathfrak{g}$} on a tensor product of unitarizable $\mathfrak{g}$}-modules.
3.1. Feigin–Frenkel centers
Let $t$ be an even variable, and let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie (super)algebra. We denote by $V_{{\rm crit}}({\mathfrak{g}})$ the vacuum module at the critical level over the central extension $\widehat{{\mathfrak{g}}}:={\mathfrak{g}}[t,t^{-1}]\oplus{\mathbb{C}}K$ of the loop algebra ${\mathfrak{g}}[t,t^{-1}]$. By the Poincaré–Birkhoff–Witt theorem, $V_{{\rm crit}}({\mathfrak{g}})$ can be identified with $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ as (super)spaces. Moreover, the $\widehat{{\mathfrak{g}}}-$module structure on $V_{{\rm crit}}({\mathfrak{g}})$ allows us to define a vertex algebra structure on $V_{{\rm crit}}({\mathfrak{g}})$, called the universal affine vertex algebra at the critical level (see [Fr, FBZ, Mo2, MR] for details).
[…]
${\mathfrak{z}}(\widehat{{\mathfrak{g}}})=\left\{\,\!v\in V_{{\rm crit}}({\mathfrak{g}})\;\middle|\;{\mathfrak{g}}[t]v=0\!\,\right\}.$
[…]
${\rm T}(1)=0\qquad\text{and}\qquad{\rm T}(A[-r])=rA[-r-1]$ (3.1)
[…]
The following result is a famous theorem of Feigin and Frenkel.
Theorem 3.1 ([FF]).
[…]
Remark 3.2.
Theorem 3.1 holds as well for the general linear Lie algebra (see [CM, Theorem 3.1] (and also [CT]).
Let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$, which has rank $d$. For ${\mathfrak{x}}=\mathfrak{a,c,d}$, recall the elements $e^{{\mathfrak{x}}}_{ij}$ defined in (2.1), (2.2) and (2.3), respectively. We denote by $\mathfrak{g}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{g}$} spanned by all the $e^{\mathfrak{x}}_{ij}$ and by $\mathfrak{h}$$_{\mathbb{R}}$ the $\mathbb{R}$}-subspace of $\mathfrak{h}$} spanned by all the $e^{\mathfrak{x}}_{ii}$.
Let $U({\mathfrak{g}}[t])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U({\mathfrak{g}}[t])$ generated by $A[r]$ for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{Z}}_{+}$, and let $U(t^{-1}{\mathfrak{g}}[t^{-1}])_{\mathbb{R}}$ be the $\mathbb{R}$}-subalgebra of $U(t^{-1}{\mathfrak{g}}[t^{-1}])$ generated by A[r]A[-r] for $A\in{\mathfrak{g}}_{\mathbb{R}}$ and $r\in{\mathbb{N}}$. The $\mathbb{R}$}-algebras $U({\mathfrak{h}}[t])_{\mathbb{R}}$ and $U(t^{-1}{\mathfrak{h}}[t^{-1}])_{\mathbb{R}}$ are defined analogously.
By [CM, Theorem 3.1] for type $\mathfrak{a}$ and [Mo3, Theorem 2.3] for type $\mathfrak{c,d}$ (see also [Mo1, Ya]), there exist
[…]
@@ 3.2. Bethe algebras
Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra. For any $\mu\in{\mathfrak{g}}^{*}$, there is an algebra homomorphism
$\Psi^{\mu}:U(t^{-1}{\mathfrak{g}}[t^{-1}])\longrightarrow U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$
@@
$\Psi^{\mu}(A[-r])=A\otimes(t-z)^{-r}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$ (3.3)
[…]
$\underline{{\rm ev}}_{\bf z}:U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]\longrightarrow U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt],$
such that $\underline{{\rm ev}}_{\bf z}(z^{-1})=z^{-1}$. We have the commutative diagram
$U(t^{-1}{\mathfrak{g}}[t^{-1}])$$U({\mathfrak{g}}[t])[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$U({\mathfrak{g}})^{\otimes\ell}[\kern-1.1pt[z^{-1}]\kern-1.1pt]$$\scriptstyle{\Psi^{\mu}_{\bf z}}$$\scriptstyle{\Psi^{\mu}}$$\scriptstyle{\underline{{\rm ev}}_{\bf z}}$
where $\Psi^{\mu}_{\bf z}$ is defined by
$\Psi^{\mu}_{\bf z}(A[-r])=\sum_{i=1}^{\ell}\frac{A^{(i)}}{(z_{i}-z)^{r}}+\delta_{r,1}\mu(A),\qquad\mbox{for $A\in{\mathfrak{g}}\,$ and $\,r\in{\mathbb{N}}$.}$
[…]
$\big(\mathcal{A}_{{\mathfrak{g}}}^{\mu}({\bf z})\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu}=\big(\mathcal{B}_{\mathfrak{g}}^{\mu}\big)_{\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})}.$
In this paper, we will restrict our attention to $\mathcal{B}_{\mathfrak{g}}^{\mu}$.
For $\mu\in{\mathfrak{g}}^{*}$, we denote by ${\mathfrak{g}}^{\mu}$ the centralizer of $\mu$ in $\mathfrak{g}$}. Moreover, $\mu$ is called regular if ${\mathfrak{g}}^{\mu}$ has dimension equal to the rank of $\mathfrak{g}$}. Set
${\bm{X}}_{\!\ell}=\left\{\,\!(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}\;\middle|\;z_{i}\not=z_{j}\,\,\mbox{for any $i\not=j$}\!\,\right\},$
called the configuration space of $\ell$ distinct points on $\mathbb{C}$$^{\ell}$.
Let $M_{1},\ldots,M_{\ell}$ be $\mathfrak{g}$}-modules, and let ${\bf z}\in{\bm{X}}_{\!\ell}$. If ${\mathfrak{h}}\subseteq{\mathfrak{g}}^{\mu}$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ commutes with $\mathfrak{h}$} and hence acts on the $\gamma$-weight space $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})_{\gamma}$ of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\gamma\in{\mathfrak{h}}^{*}$ ([Ry1, Proposition 4]).
Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules. We have the following remarkable results about the action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$.
Theorem 3.3 ([FFRy, Corollary 5]).
@@
The space $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\mu}$-module for any regular $\mu\in{\mathfrak{g}}^{*}$ and ${\bf z}\in{\bm{X}}_{\!\ell}$.
Theorem 3.4 ([FFRy, Corollary 6]).